OpenStudy (anonymous):

Renee took a long flight to visit her family that lives out of the country. She slept for the first quarter of the trip and watched movies for half of the remaining flight. If the total trip took nine and three-fourths hours, how long did Renee watch movies

OpenStudy (anonymous):

@LukeBlueFive can you help please

OpenStudy (anonymous):

@KamiBug can you please help

OpenStudy (kamibug):

She watched movies for 4.75 hours. (Or 4 3/4)

OpenStudy (lukebluefive):

Write out the information in equation form: \[sleep = \frac{1}{4} * trip\] \[movies = \frac{1}{2}*(trip - sleep)\] Plug in 9.75 for trip: \[sleep = \frac{1}{4} * 9.75\] Then plug that result into the movie equation: \[movies = \frac{1}{2}*(9.75- sleep)\]